sadie5:

Which equation is true for the value b = 10? A. 2(b + 4) = 16 B. 2(b + 2) = 40 C. 3(b – 2) = 24 D. 2(8 + b) = 42 E. 3(b – 4) = 20

563blackghost:

The easiest way to solve the question is plugging in `10` for each variable of `b`. \(\large\bf{A. 2 (10+4) = 16 ?}\) \(\large\bf{B. 2(10+2)=40 ?}\) \(\large\bf{C. 3(10-2)=24?}\) \(\large\bf{D. 2(8+10)=42?}\) \(\large\bf{E. 3(10-4)=20?}\)

563blackghost:

Simplify each to see if the equation stands true.

KynxKatz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @563blackghost Simplify each to see if the equation stands true. \(\color{#0cbb34}{\text{End of Quote}}\) Can you tell hiim how you got 10?

KynxKatz:

Oh nvm

supie:

Yes, you can do that ways you can solve the equation is solve by using the distributive property, where you multiply what is outside of the parenthesis by everything that is inside of the parenthesis, \(a(b+c)=ab+ac\) or you can solve what is inside of the parenthesis then multiply that by what is outside of the parenthesis then whichever one is true would be your answer